### v2

@@ G.5 N-bus Lossy Power System
Unlike N-bus lossless power systems in Appendix G.3 which has a well-known energy-based storage function served as a valid Lyapunov function for asymptotical stability guarantee, the N-bus lossy power system (Cui & Zhang, 2022) does not have a known analytical Lyapunov function to certify stability, though by passivity the system should be asymptotically stable at origin. Utilizing proposed framework, we aim to discover a valid analytical Lyapunov function for N-bus lossy power system to formally certify its asymptotically stability.
The $\theta_{i}$ and $\delta_{i}$ are the angle and frequency deviation of bus $i$, the dynamics of N-bus lossy power system is represented by the swing equation, formulated as:
@@
where $m_{i}$ is the generator inertia constant, $d_{i}$ is the combined frequency response coefficient from synchronous generators and frequency sensitive load, and $p_{i}$ is the net power injection, for each bus $i=1,\cdots,N$. The susceptance and conductance of the line $(i,j)$ are $B_{ij}=B_{ji}$ and $G_{ij}=G_{ji}$, respectively; and 0 if the buses are not connected. Because power systems do not have real-time communication infrastructure, we restrict input $u_{i}$ to be a static feedback controller where only its local frequency measurement $\omega_{i}$ is available. Like lossless power system, since the frequency dynamics of the system depends only on the phase angle differences, so we change the coordinates:
@@
We test proposed framework on 2-bus lossy power system. In experiment, we set $p_{i}=1$,$m_{i}=2$, $d_{i}=1$, $u_{i}(w_{i})=w_{i}$, $B_{ij}=1,G_{ij}=1\;\forall\;i\neq j$, $B_{ii}=0,G_{ij}=0$. By this setting, the equilibrium point for this system is at origin, i.e. $\delta_{i}^{*}=\omega_{i}^{*}=0$. The state space for the experiment is defined as: $\mathcal{D}=\{(\delta_{1},\delta_{2},\omega_{1},\omega_{2})\in\mathbb{R}^{4}\bm{\mid}|\delta_{i}|\leq 0.75\text{ and }|\omega_{i}|\leq 2\text{ for }i=1,2\}$. The proposed method found two valid Lyapunov function $V(\delta_{1},\delta_{2},\omega_{1},\omega_{2})=\omega_{1}^{2}+\omega_{2}^{2}+(\omega_{2}-\sin(\delta_{1})+\sin(\delta_{2}))^{2}$ and $V(\delta_{1},\delta_{2},\omega_{1},\omega_{2})=\omega_{1}^{2}+\omega_{2}^{2}+(-\omega_{1}-\sin(\delta_{1})+\sin(\delta_{2}))^{2}$. Both Lyapunov functions pass the formal verification by SMT solver in the state space $\mathcal{D}\backslash\mathcal{B}_{\epsilon}(0)$, where precision $\delta$ is set to be $e^{-12}$ and $\epsilon=e^{-3}$ to avoid tolerable numerical error.
@@ H.1 Ablation Study I - Risk-seeking Policy Gradient
To better understand the influence of $\alpha$ in policy gradient to the training performance of the symbolic transformer, we experiment our framework without GP refinement on 3-D Trig dynamics under different choices of $\alpha$ (the risk-seeking quantile) over 100 epochs. Specifically, we choose $\alpha=0.1,0.5,1$ and compare the transformer model’s performance by the reward trajectories and success discovery rate. Figure 5 plots the reward trajectories—both the highest reward and the $90\%$-quantile—under the three different choices of $\alpha$ (the risk-seeking quantile). For $\alpha=0.1$, the setting we used on tested dynamics in experiments, the highest reward and the 90% quantile converge steadily to 1 and 0.9, respectively, exhibiting lower variance and faster stabilization. For $\alpha=0.5$, the median-based threshold increases steadily but shows greater instability. Finally, for $\alpha=1$ vanilla policy gradient, the best reward reaches only 0.94, while the 90% quantile hovers around 0.6. Additionally, with $\alpha=0.1$, the framework achieves the highest success recovery rate ($66.67\%$) compared to $33.33\%$ and $0\%$ for $\alpha=0.5$ and $\alpha=1$. These findings correspond to the analysis of misaligned objective between standard policy gradient and Lyapunov function construction in Subsection 4.3, confirm the importance of risk-seeking policy optimization, and highlight the effectiveness of setting $\alpha=0.1$ in experiments.
@@ H.2 Ablation Study II - Global-optimization-based Numerical Verification Process
This work proposes a global-optimization-based numerical verification for candidate verification and counter-examples’ feedback within the training paradigm. To identify the effectiveness of proposed verification algorithm, this ablation study compares the following verification methods on proposed framework without GP component on the 6-D polynomial dynamics (Appendix F.4):
@@
  • Root finding: Proposed framework with root-finding numerical verification introduced in Feng et al. (2024).
@@
Table 4: Number of epoch, training time, success rate, and formal verification time comparison between four different settings in Ablation study II. Runtime is the average training time for successful trials. The Succ. % is the success rate of finding a valid Lyapunov function out of 3 random seeds. Ver. Time is the average formal verification time of identified candidates on successful trials. The experiment terminates either a valid Lyapunov function is found (pass formal verification), the training epoch exceeds 250, or timeout (4 hours limit).
@@
Figure 7 compares the highest reward trajectory in each epoch for all four settings, and Table 4 summarizes the training and formal verification statistics. SMT is not included in Figure 7 as the training stuck on the SMT solver for hours in the middle of the training stage. The proposed SHGO-based counterexample feedback makes the training reward an adversarial reward, as it can identify the most critical violations. As a ‘challenger’, the global optimization method finds the adversarial counterexamples, not just mild violations. Consequently, this leads to a faster refinement of the policy to mitigate the critical violations, and results in stronger guarantees on the final Lyapunov candidates. However, it risks training stability and may over-focus on ‘hard’ violations. In this case, it identifies a valid Lyapunov function $V_{1}(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$ which is well-structured and easy to verify by formal verification tools. Note that, unlike in the experiment section, the GP component is not used here, resulting in a less structured local Lyapunov function with a few squared interaction terms $(x_{i}+x_{j})^{2}$ under one random seed. This increases verification time compared to Table 1, yet remains significantly faster than verifying the final result from the random sampling scheme.
@@
Root-finding usually identifies counterexamples near function roots, producing mostly mild violations. Although an ideal root-finding algorithm might be efficient, off-the-shelf methods become inaccurate for dimensions above five. As a result, they fail to provide precise violation signals or strong optimization guidance to the transformer, causing the best reward to stall at $0.7$. Similarly, dreal SMT often locates only mild violations rather than critical ones, limiting its impact on optimization. While these counterexamples help the transformer reduce violations, the refined transformer increasingly acts as an adversary to the SMT solver—prolonging verification until timeout. In neural network + formal verification settings, this effect is even exacerbated by the model’s overparameterization. Consequently, we rely on global optimization instead of formal verification in the training loop.
@@ H.3 Ablation Study III - Generic Programming and Expert Guidance Loss
  • Transformer + GP: A symbolic transformer model with GP component for candidate refinement but without expert-guided learning.
  • Transformer + GP + expert-guided learning (ours): A symbolic transformer model with GP component for both candidate refinement and expert-guided learning for transformer optimization.
@@
Figure 8: The figure on the left visualizes the overall best reward (best candidate among $\tilde{\mathcal{V}}_{\phi}\cup\tilde{\mathcal{V}}_{gp}$) trajectory. The figure on the left visualizes the best reward trajectory for transformer candidates (best candidate among $\tilde{\mathcal{V}}_{\phi}$ only, excludes GP-generated solutions). This two plots compare the performance of transformer + GP refinement and ours method in Ablation Study III.

### v3

@@ G.5 N-bus Lossy Power System
Unlike N-bus lossless power systems in Appendix G.3 which has a well-known energy-based storage function served as a valid Lyapunov function for asymptotical stability guarantee, the N-bus lossy power system (Cui & Zhang, 2022) does not have a known analytical Lyapunov function to certify stability, though by passivity the system should be asymptotically stable at origin. Utilizing the proposed framework, we aim to discover a valid analytical Lyapunov function for an N-bus lossy power system to formally certify its asymptotic stability.
The $\theta_{i}$ and $\delta_{i}$ are the angle and frequency deviation of bus $i$, the dynamics of an N-bus lossy power system is represented by the swing equation, formulated as:
@@
where $m_{i}$ is the generator inertia constant, $d_{i}$ is the combined frequency response coefficient from synchronous generators and frequency-sensitive load, and $p_{i}$ is the net power injection, for each bus $i=1,\cdots,N$. The susceptance and conductance of the line $(i,j)$ are $B_{ij}=B_{ji}$ and $G_{ij}=G_{ji}$, respectively. The value is 0 if the buses are not connected. In this work, we consider input $u_{i}$ to be a static feedback controller where only its local frequency measurement $\omega_{i}$ is available. Like the lossless power system, since the frequency dynamics of the system depends only on the phase angle differences, we change the coordinates:
@@
We test the proposed framework on a 2-bus lossy power system. In experiment, we set $p_{i}=1$,$m_{i}=2$, $d_{i}=1$, $u_{i}(w_{i})=w_{i}$, $B_{ij}=1,G_{ij}=1\;\forall\;i\neq j$, $B_{ii}=0,G_{ij}=0$. By this setting, the equilibrium point for this system is at the origin, i.e. $\delta_{i}^{*}=\omega_{i}^{*}=0$. The state space for the experiment is defined as: $\mathcal{D}=\{(\delta_{1},\delta_{2},\omega_{1},\omega_{2})\in\mathbb{R}^{4}\bm{\mid}|\delta_{i}|\leq 0.75\text{ and }|\omega_{i}|\leq 2\text{ for }i=1,2\}$. The proposed method found two valid Lyapunov function $V(\delta_{1},\delta_{2},\omega_{1},\omega_{2})=\omega_{1}^{2}+\omega_{2}^{2}+(\omega_{2}-\sin(\delta_{1})+\sin(\delta_{2}))^{2}$ and $V(\delta_{1},\delta_{2},\omega_{1},\omega_{2})=\omega_{1}^{2}+\omega_{2}^{2}+(-\omega_{1}-\sin(\delta_{1})+\sin(\delta_{2}))^{2}$. Both Lyapunov functions pass the formal verification by SMT solver in the state space $\mathcal{D}\backslash\mathcal{B}_{\epsilon}(0)$, where precision $\delta$ is set to be $e^{-12}$ and $\epsilon=e^{-3}$ to avoid tolerable numerical error.
G.6 9-D Synthetic Dynamics
Consider the synthetic dynamics adapted from Appendices F.4 & G.2 with linear interactions between two subsystems:
$\dot{x}_{1}$ $=-x_{1}+0.5x_{2}-0.1x_{5}^{2},$
$\dot{x}_{2}$ $=-0.5x_{1}-x_{2}+0.1x_{8},$
$\dot{x}_{3}$ $=-x_{3}+0.5x_{4}-0.1x_{1}^{2},$
$\dot{x}_{4}$ $=-0.5x_{3}-x_{4},$
$\dot{x}_{5}$ $=-x_{5}+0.5x_{6},$
$\dot{x}_{6}$ $=-0.5x_{5}-x_{6}+0.1x_{2}^{2},$
$\dot{x}_{7}$ $=x_{8},$
$\dot{x}_{8}$ $=-\sin(x_{7})\cos(x_{7})-x_{8}-\sin(x_{9})\cos(x_{9})-0.1x_{2},$
$\dot{x}_{9}$ $=x_{8}-x_{9}.$
To properly address the trigonometric terms in $\dot{x}_{8}$, the Lyapunov function for this dynamics can’t be a simple form like $\sum_{i=1}^{n}x_{i}^{2}$ and should include some trigonometric terms. Setting the state space $\mathcal{D}=\{x\in\mathbb{R}^{9}||x_{i}|\leq 1.5,\forall\;i=1,\cdots,9\}$, our method successfully identifies a valid Lyapunov function $V=\sum_{i=1}^{6}x_{i}^{2}+\sin(x_{7})^{2}+x_{8}^{2}-\cos(x_{9})+1$, which passes formal verification following settings in Section 5.
@@ H.1 Ablation Study I - Risk-seeking Policy Gradient
To better understand the influence of $\alpha$ in policy gradient on the training performance of the symbolic transformer, we experiment with our framework without GP refinement on 3-D Trig dynamics under different choices of $\alpha$ (the risk-seeking quantile) over 100 epochs. Specifically, we choose $\alpha=0.1,0.5,1$ and compare the transformer model’s performance by the reward trajectories and success discovery rate. Figure 5 plots the reward trajectories—both the highest reward and the $90\%$-quantile—under the three different choices of $\alpha$ (the risk-seeking quantile). For $\alpha=0.1$, the setting we used on tested dynamics in experiments, the highest reward and the 90% quantile converge steadily to 1 and 0.9, respectively, exhibiting lower variance and faster stabilization. For $\alpha=0.5$, the median-based threshold increases steadily but shows greater instability. Finally, for $\alpha=1$ vanilla policy gradient, the best reward reaches only 0.94, while the 90% quantile hovers around 0.6. Additionally, with $\alpha=0.1$, the framework achieves the highest success recovery rate ($66.67\%$) compared to $33.33\%$ and $0\%$ for $\alpha=0.5$ and $\alpha=1$. These findings correspond to the analysis of misaligned objective between the standard policy gradient and Lyapunov function construction in Subsection 4.3, confirm the importance of risk-seeking policy optimization, and highlight the effectiveness of setting $\alpha=0.1$ in experiments.
@@ H.2 Ablation Study II - Global-optimization-based Numerical Verification Process
This work proposes a global-optimization-based numerical verification for candidate verification and counterexamples’ feedback within the training paradigm. To identify the effectiveness of the proposed verification algorithm, this ablation study compares the following verification methods on the proposed framework without the GP component on the 6-D polynomial dynamics (Appendix F.4):
@@
  • Root finding: Proposed framework with root-finding numerical verification introduced in Feng et al. (2024c).
@@
Table 7: Number of epoch, training time, success rate, and formal verification time comparison between four different settings in Ablation study II. Runtime is the average training time for successful trials. The Succ. % is the success rate of finding a valid Lyapunov function out of 3 random seeds. Ver. Time is the average formal verification time of identified candidates on successful trials. The experiment terminates either when a valid Lyapunov function is found (passes formal verification), the training epoch exceeds 250, or a timeout (4 hours limit).
@@
Figure 7 compares the highest reward trajectory in each epoch for all four settings, and Table 7 summarizes the training and formal verification statistics. SMT is not included in Figure 7 as the training stuck on the SMT solver for hours in the middle of the training stage. The proposed SHGO-based counterexample feedback makes the training reward an adversarial reward, as it can identify the most critical violations. As a ‘challenger’, the global optimization method finds the adversarial counterexamples, not just mild violations. Consequently, this leads to a faster refinement of the policy to mitigate the critical violations and results in stronger guarantees on the final Lyapunov candidates. However, it risks training stability and may over-focus on ‘hard’ violations. In this case, it identifies a valid Lyapunov function $V_{1}(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$ which is well-structured and easy to verify by formal verification tools. Note that, unlike in the experiment section, the GP component is not used here, resulting in a less structured local Lyapunov function with a few squared interaction terms $(x_{i}+x_{j})^{2}$ under one random seed. This increases verification time compared to Table 1, yet remains significantly faster than verifying the final result from the random sampling scheme.
@@
Root-finding usually identifies counterexamples near function roots, producing mostly mild violations. Although an ideal root-finding algorithm might be efficient, off-the-shelf methods become inaccurate for dimensions above five. As a result, they fail to provide precise violation signals or strong optimization guidance to the transformer, causing the best reward to stall at $0.7$. Similarly, dreal SMT often locates only mild violations rather than critical ones, limiting its impact on optimization. While these counterexamples help the transformer reduce violations, the refined transformer increasingly acts as an adversary to the SMT solver, prolonging verification until timeout. In neural network + formal verification settings, this effect is even exacerbated by the model’s overparameterization. Consequently, we rely on global optimization instead of formal verification in the training loop.
@@ H.3 Ablation Study III - Generic Programming and Expert Guidance Loss
  • Transformer + GP: A symbolic transformer model with a GP component for candidate refinement but without expert-guided learning.
  • Transformer + GP + expert-guided learning (ours): A symbolic transformer model with a GP component for both candidate refinement and expert-guided learning for transformer optimization.
@@
Figure 8: The figure on the left visualizes the overall best reward (best candidate among $\tilde{\mathcal{V}}_{\phi}\cup\tilde{\mathcal{V}}_{gp}$) trajectory. The figure on the left visualizes the best reward trajectory for transformer candidates (best candidate among $\tilde{\mathcal{V}}_{\phi}$ only, excludes GP-generated solutions). These two plots compare the performance of transformer + GP refinement and our method in Ablation Study III.